fj02114:

Find the value of x in the expression 2003-02-05-00-00_files/i0080000.jpg.

Vocaloid:



Vocaloid:

remember your exponent rules (a^b)^d = a^(b*d) applying this logic to your problem gives us 2^x = 4 (a^4)^x = a^8 (b^2)^x = b^4 based on this, any ideas what x might be?